In a study of breast cancer, a mother was considered exposed if she first gave birth at 30 years of age or older. Among 1,628 mothers who were exposed, 31 subsequently developed breast cancer. Among the 4,450 mothers who first gave birth before the age of 30, 65 subsequently developed breast cancer. What is the appropriate conclusion regarding the attributable risk of breast cancer due to an older age at first birth (per 1,000)?

Across the US, give birth for the first time at 30 years of age or older causes 440 cases of breast cancer

Among 1,000 mothers who first gave birth at 30 years of age or older, 4.4 cases of breast cancer can be attributed to an older age at first birth

Eliminating births to women older than 30 would prevent 4.4% of cases of breast cancer

4.4 cases of breast cancer are attributed to women giving birth for the first time at 30 years of age or older among every 1,000 mothers in the population

None of the answers listed
Answer #1

ANSWER- NONE OF THE ANSWERS LISTED

( As in first case where she first gave birth at 30 years of age or older p= 31/ 1628 =0.01904 i.e 19.4 breast cancer for 1000

And in the second case where she first gave birth before the age of 30, p= 64/4450 = 0.01438 = 14.38 breast cancer for 1000)